The width of a rectangle is 10 inches longer than its length. If the area of the rectangle is 56 square inches , which equation could be used to determine its length?
Answer by Boreal(15235)

length=x
width=x+10
x*(x+10)=56
x^2+10x-56=0. That can find the length.
(x+14)(x-4)=0
x=4 inches length and x+10=14 inches width.